# The structure of the $f_{0}(980)$ from system size dependent hadronic resonance ratios in p+p, p+Pb, and Pb+Pb collisions at the LHC

### v1

@@ I Introduction
The increasing number of known hadronic states Workman et al. (2022a) has led to the establishment of the quark model Gell-Mann (1964) describing color-neutral hadrons as bound multi quark states. Although the internal structure of most of the known particles can be well described by the quark model as $|\bar{q}q\rangle$ (meson) or $|qqq\rangle$ (baryon) states, the internal structure of some particles, especially of hadronic resonances like the scalar $f$ mesons, remains unclear Jaffe (1977); Maiani et al. (2004); Cho et al. (2011). Alternative internal structures which are allowed by the quark model are tetra- or penta-quark states, glueballs, molecular meson+meson states or hybrids. Most prominent among these mysterious hadrons is the $f_{0}(980)$ state which is difficult to describe within the conventional quark model Achasov and Ivanchenko (1989). It has been proposed as a $|\bar{u}u+\bar{d}d\rangle$ state Chen (2003), a $|\bar{K}K\rangle$ state Weinstein and Isgur (1990); Ahmed and Xiao (2020), a tetra-quark state $|\bar{q}\bar{q}qq\rangle$ Achasov and Ivanchenko (1989); Weinstein and Isgur (1983); Achasov et al. (2021); Zhao et al. (2022), or a quark-antiquark gluon hybrid Birse et al. (1996). However, final scrutinizing evidence remains to be found. Experimentally, measurements by BESIII have observed it as a decay product and investigated its mixing with the $a_{0}(980)$ Ablikim et al. (2012, 2018), which itself may contain tetra-quark contribution Bulava et al. (2023). The study of light exotics such as the $f_{0}(980)$ was also identified as one of the major goals of the PANDA experiment Barucca et al. (2021).
@@
Thus, we describe $f_{0}(980)$ production in p+p and Pb+Pb collisions at the LHC employing a) the partial chemical equilibrium hadron resonance gas model within Thermal-FIST and b) the hybrid UrQMD model to investigate $f_{0}(980)$ production based on different assumptions of its quark structure and decay channels.
@@ II.1.1 Canonical suppression
[…]
@@ II.1.2 Partial chemical equilibrium
Figure 1: (Color online) Parameterized chemical (solid blue line) and kinetic freeze-out (dashed blue line) temperatures providing the best fits to Pb+Pb data at $\sqrt{s_{\mathrm{NN}}}=2.76$ TeV Vovchenko and Stoecker (2019); Motornenko et al. (2020a) and used in the partial chemical equilibrium thermal model (Thermal-FIST) in this work. The particilization (Cooper-Frye) temperature of the hybrid UrQMD simulation is shown as a red line.
@@ II.1.3 Other implementation details
We use an open-source package Thermal-FIST (version 1.4.2) Vovchenko and Stoecker (2019); Vovchenko (2018) in all our thermal model-based calculations. The charged multiplicity dependence of $T_{\rm ch}$, $T_{\rm kin}$, $V_{c}$, and, where applicable, $\gamma_{S}$, is described in the previous subsections. One should note that Thermal-FIST presently implements PCE in the grand-canonical ensemble only, thus, the simultaneous PCE+CSM calculation proceeds in two steps. First, we calculate hadron abundances within the chemical equilibrium CSM. Second, we use the PCE model in the grand canonical ensemble to calculate the correction factors for the yields of short-lived resonances due to hadronic phase. The abundances computed in step one are then multiplied by these factors. In practice, the PCE effects are mainly relevant in large systems, $\langle dN_{\rm ch}/d\eta\rangle\gtrsim 100$ Motornenko et al. (2020a), while canonical suppression kicks in in small systems, $\langle dN_{\rm ch}/d\eta\rangle\lesssim 100$ Vovchenko and Stoecker (2019) for hadrons under consideration in the present work. Thus, the canonical and PCE effects can be largely considered independently.
@@ II.3 Known resonance ratios
To benchmark the hybrid UrQMD model and the Thermal-FIST model with PCE+CSM, and validate their ability to describe other resonance multiplicities, we first calculate the centrality dependence of well-studied resonance ratios, i.e. the $K^{*0}(892)/K^{-}$, $\Delta^{++}(1232)/p$ and $\rho^{0}(770)/\pi$ ratios, as function of system size. For the sake of brevity we will drop the round brackets depicting the mass number. Fig. 2 shows the centrality dependence of the $K^{*0}/K^{-}$ (left panel), $\Delta^{++}/p$ (middle panel) and $\rho^{0}/\pi$ (right panel) ratios compared to ALICE data Abelev et al. (2015); Acharya et al. (2022a, 2023b, b, 2019). The hybrid UrQMD calculations are shown for Pb+Pb collisions with hadronic phase (full red circles) and without hadronic phase (solid black line), for p+p collisions (full red squares) while the Thermal-FIST calculations with PCE+CSM are shown as a dashed blue line and the experimental data is depicted as green symbols.
Figure 2: (Color online) The centrality dependence of the $K^{*0}/K^{-}$ (left panel), $\Delta^{++}/p$ (middle panel) and $\rho^{0}/\pi$ (right panel) ratios compared to ALICE data from different system sizes. The hybrid UrQMD calculations are shown for Pb+Pb collisions with hadronic phase (full red circles) and without hadronic phase (solid black line), for p+p collisions (full red squares) while the Thermal-FIST calculations with PCE+CSM are shown as a dashed blue line and the experimental data is depicted as green symbols Abelev et al. (2015); Acharya et al. (2022a, 2023b, b, 2019); Adam et al. (2016); Acharya et al. (2020, 2019).
@@ III Results
ALICE has measured the $f_{0}(980)$ recently in the invariant mass distribution of $\pi^{+}\pi^{-}$ pairs in the invariant mass range from $0.8\leq M_{\pi^{+}\pi^{-}}\leq 1.6$ GeV Acharya et al. (2023a). The 2023 PDG report Workman et al. (2022b) still quotes the $f_{0}(980)$ decay channels into $\pi\pi$ and $\overline{K}K$ as “seen” leaving the branching ratios unknown as well [^4] Note, that ALICE assumes a $30\%$ branching into $\overline{K}K$.. This leaves us the freedom to explore dependences of the observable multiplicity on the branching ratio into kaons. To do so, we employ two different branching ratios: a) a 0% branching ratio into kaons (only pion decay) and b) a 30% branching ratio into kaons and reduced branching to pions. Fig. 4 shows the invariant mass distribution of the reconstructable $f_{0}(980)$ resonances reconstructed in the pion channel from central Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV from UrQMD. The solid line shows the invariant mass distributions for UrQMD calculations with a 30% branching ratio for the kaon decay branch of the $f_{0}(980)$ while the calculations shown by the dashed line assume a 0% branching ratio to the kaon channel. The vertical line shows the lower reconstruction limit reported by ALICE.
@@ IV Discussion
In general, the estimated yield of observable $f_{0}(980)$ resonances is significantly lower in all models compared to the ALICE p+p data. This and the above conclusions make additional measurements of the $f_{0}$ in peripheral Pb+Pb collisions with ALICE, as well as future measurements at lower beam energies, for instance with the CBM experiment, necessary and worthwhile.

### v2

@@ I Introduction
The increasing number of known hadronic states Workman et al. (2022a) has led to the establishment of the quark model Gell-Mann (1964) describing color-neutral hadrons as bound multi quark states. Although the internal structure of most of the known particles can be well described by the quark model as $|\bar{q}q\rangle$ (meson) or $|qqq\rangle$ (baryon) states, the internal structure of some particles, especially of hadronic resonances like the scalar $f$ mesons, remains unclear Jaffe (1977); Montanet (1983); Close (1988); Morgan and Pennington (1991); Tornqvist (1982); Tornqvist and Roos (1996); Morgan (1974); Maiani et al. (2004); Cho et al. (2011). Alternative internal structures which are allowed by the quark model are tetra- or penta-quark states, glueballs, molecular meson+meson states or hybrids. Most prominent among these mysterious hadrons is the $f_{0}(980)$ state which is difficult to describe within the conventional quark model Achasov and Ivanchenko (1989). It has been proposed as a $|\bar{u}u+\bar{d}d\rangle$ state Chen (2003), a $|\bar{K}K\rangle$ state Weinstein and Isgur (1990); Ahmed and Xiao (2020); Wang et al. (2022), a tetra-quark state $|\bar{q}\bar{q}qq\rangle$ Achasov and Ivanchenko (1989); Weinstein and Isgur (1983); Achasov et al. (2021); Zhao et al. (2022), or a quark-antiquark gluon hybrid Birse et al. (1996). See also Oller et al. (1998, 1999); Guerrero and Oller (1999) for advances on the description of scalar resonances in the two meson scattering states.
In addition, only few attempts at understanding the modifications of the $f_{0}(980)$ in dense matter exist (see e.g. Oset and Vicente Vacas (2000)), mostly due to its unclear nature. Nevertheless, a better understanding of the $f_{0}(980)$ may help to shed light on the nature of (the restoration of) chiral symmetry in hot and dense matter, if it could be better understood.
However, final scrutinizing evidence remains to be found. Experimentally, measurements by BESIII have observed it as a decay product and investigated its mixing with the $a_{0}(980)$ Ablikim et al. (2012, 2018), which itself may contain tetra-quark contribution Bulava et al. (2023). The study of light exotics such as the $f_{0}(980)$ was also identified as one of the major goals of the PANDA experiment Barucca et al. (2021).
@@
Thus, we describe $f_{0}(980)$ production in pp and PbPb collisions at the LHC employing a) the partial chemical equilibrium hadron resonance gas model within Thermal-FIST and b) the hybrid UrQMD model to investigate $f_{0}(980)$ production based on different assumptions of its quark structure and decay channels.
Here, we have to note that in both, the UrQMD model and Thermal-FIST, the $f_{0}$ is treated as an independent quasiparticle degree of freedom. This means that the actual quark structure does not enter explicitly in either approach but implicitly through its production mechanism (e.g. as an $s+\overline{s}$ pair in the string or by including a strangeness suppression factor in Thermal-FIST) or its decays (either to $\pi+\pi$ or $K+\overline{K}$). More exotic quark configurations would therefore be difficult to implement.
@@ II.1.1 Canonical suppression
The correlation volume $V_{c}$ corresponds to three units of rapidity for all $\langle dN_{\rm ch}/d\eta\rangle$, i.e. $V_{c}=3dV/dy$. The volume per rapidity $dV/dy$ itself is a linear function of multiplicity, $dV/dy=v\langle dN_{\rm ch}/d\eta\rangle$, where $v=2.4$ fm3.
@@ II.1.2 Partial chemical equilibrium
Figure 1: Parameterized chemical (solid blue line) and kinetic freeze-out (dashed blue line) temperatures providing the best fits to Pb+Pb data at $\sqrt{s_{\mathrm{NN}}}=2.76$ TeV Vovchenko and Stoecker (2019); Motornenko et al. (2020a) and used in the partial chemical equilibrium thermal model (Thermal-FIST) in this work. The particlization (Cooper-Frye) temperature of the hybrid UrQMD simulation is shown as a red line.
@@ II.1.3 Other implementation details
We use an open-source package Thermal-FIST (version 1.4.2) Vovchenko and Stoecker (2019); Vovchenko (2018) in all our thermal model-based calculations. The charged multiplicity dependence of $T_{\rm ch}$, $T_{\rm kin}$, $V_{c}$, and, where applicable, $\gamma_{S}$, is described in the previous subsections. One should note that Thermal-FIST presently implements PCE in the grand-canonical ensemble only, thus, the simultaneous PCE+CSM calculation proceeds in two steps. First, we calculate hadron abundances within the chemical equilibrium CSM. Second, we use the PCE model in the grand canonical ensemble to calculate the correction factors for the yields of short-lived resonances due to hadronic phase. The abundances computed in step one are then multiplied by these factors. In practice, the PCE effects are mainly relevant in large systems, $\langle dN_{\rm ch}/d\eta\rangle\gtrsim 100$ Motornenko et al. (2020a), while canonical suppression becomes relevant in small systems, $\langle dN_{\rm ch}/d\eta\rangle\lesssim 100$ Vovchenko and Stoecker (2019) for hadrons under consideration in the present work. Thus, the canonical and PCE effects can be largely considered independently.
@@
Figure 2: (Color online) The centrality dependence of the $K^{*0}/K^{-}$ (left panel), $\Delta^{++}/p$ (middle panel) and $\rho^{0}/\pi$ (right panel) ratios compared to ALICE data from different system sizes. The hybrid UrQMD calculations are shown for Pb+Pb collisions, at $\sqrt{s_{NN}}=5.02$ TeV, with a hadronic phase (full red circles) and without hadronic phase (solid black line) and for p+p collisions (full red squares). The Thermal-FIST calculations with PCE+CSM are shown as a dashed blue line. The experimental data for pp (dark yellow), pPb (purple) and PbPb (green) collisions at $\sqrt{s_{NN}}=5.02$ (triangles) and $2.7$ TeV (diamonds) are also shown Abelev et al. (2015); Acharya et al. (2022a, 2023b, b, 2019); Adam et al. (2016); Acharya et al. (2020).
@@ II.3 Known resonance ratios
To benchmark the hybrid UrQMD model and the Thermal-FIST model with PCE+CSM, and validate their ability to describe other resonance multiplicities, we first calculate the centrality dependence of well-studied resonance ratios, i.e. the $K^{*0}(892)/K^{-}$, $\Delta^{++}(1232)/p$ and $\rho^{0}(770)/\pi$ ratios, as function of system size. For the sake of brevity we will drop the round brackets depicting the mass number. Fig. 2 shows the centrality dependence of the $K^{*0}/K^{-}$ (left panel), $\Delta^{++}/p$ (middle panel) and $\rho^{0}/\pi$ (right panel) ratios compared to ALICE data Abelev et al. (2015); Acharya et al. (2022a, 2023b, b, 2019); Adam et al. (2016); Acharya et al. (2020). The hybrid UrQMD calculations are shown for Pb+Pb collisions with hadronic phase (full red circles) and without hadronic phase (solid black line), for p+p collisions (full red squares) while the Thermal-FIST calculations with PCE+CSM are shown as a dashed blue line and the experimental data is depicted as green symbols.
@@ III Results
For our current study, there are several reasons why we only discuss the hadronic decay channel. The most important one is that the whole conclusion of our work relies on the effect of the hadronic rescattering phase of the collision and how it affects the measurable $f_{0}$. Those $f_{0}$ which decay in the di-electron channel are lost for any rescattering and thus carry different information. In addition, the electromagnetic branching fraction of the $f_{0}$ is not well known and would simply introduce more uncertainty. Finally, the measurement of the $f_{0}$ in the di-lepton channel is very challenging due to the background of other vector meson decays as well as q-qbar contributions.
@@
ALICE has measured the $f_{0}(980)$ recently in the invariant mass distribution of $\pi^{+}\pi^{-}$ pairs in the invariant mass range from $0.8\leq M_{\pi^{+}\pi^{-}}\leq 1.6$ GeV Acharya et al. (2023a). The cut in the invariant mass range for the experimental reconstruction is necessary to reduce the background of other resonances including the other $f_{0}$ states as well as the $\rho$ meson. Since our simulation allows us to exclude such background we have the full mass range available. The 2023 PDG report Workman et al. (2022b) still quotes the $f_{0}(980)$ decay channels into $\pi\pi$ and $\overline{K}K$ as “seen” leaving the branching ratios unknown as well [^4] Note, that ALICE assumes a $30\%$ branching into $\overline{K}K$.. This leaves us the freedom to explore dependences of the observable multiplicity on the branching ratio into kaons. To do so, we employ two different branching ratios: a) a 0% branching ratio into kaons (only pion decay) and b) a 30% branching ratio into kaons and reduced branching to pions.  [^5] In principle the branching ratio to the $\overline{K}K$ channel may even be larger than $30\%$ Ahmed and Xiao (2020); Wang et al. (2022). However, as we will see, in that case the results would simply resemble those without a hadronic rescattering, i.e. already at $30\%$ the effect of the rescattering is very minor. This is also seen in the Thermal-FIST calculation which either employs a $100\%$ pion decay OR a $100\%$ $\overline{K}K$ decay channel. To determine whether the $\overline{K}K$ is larger than $30\%$ would therefore require a very precise measurement of the centrality dependence.
Fig. 4 shows the invariant mass distribution of the reconstructable $f_{0}(980)$ resonances reconstructed in the pion channel from central Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV from UrQMD. The solid line shows the invariant mass distributions for UrQMD calculations with a 30% branching ratio for the kaon decay branch of the $f_{0}(980)$ while the calculations shown by the dashed line assume a 0% branching ratio to the kaon channel. The vertical line shows the lower reconstruction limit reported by ALICE.
@@ IV Discussion
In general, the estimated yield of observable $f_{0}(980)$ resonances is significantly lower in all models compared to the ALICE pp data, i.e. both models underestimate the $f_{0}/\pi$ ratio in pp collisions, although the errors are large. As mentioned above, an important uncertainty is the branching fraction into the two pion channel which may be significant. Furthermore, particle production in pp collisions is not necessarily perfectly described in either the UrQMD or the Thermal-FIST model. Our predictions are more reliable for larger pPb or PbPb collisions. A measurement of the $f_{0}$ in heavy-ion reactions would therefore immediately shed some light on the uncertainties in the pp results. This and the above conclusions make additional measurements of the $f_{0}$ in peripheral Pb+Pb collisions with ALICE, as well as future measurements at lower beam energies, for instance with the CBM experiment, necessary and worthwhile.